OpenStudy (anonymous):

Please help!!!!! A 6.10m × 7.60m × 2.90m room contains air at 20∘C. what is the rooms thermal energy?

OpenStudy (anonymous):

You can solve for average thermal energy of the room by using this formula KEav=(3/2)kT where k=1.38*10-23 and T= temperature in kelvin in your case 20+273=293k however i am not sure how the volume of the room is related to the changing energy?

OpenStudy (anonymous):

That's for a single molecule, I believe - a single atom, really, so if it's air you might want to add an extra degree of freedom, making it \[\bar{K}=\frac{5}{2}kT\] so if you assume it's an ideal gas, you can solve for the number of molecules based on \[PV=NkT\] \[N=\frac{PV}{kT}\] where \[P=1atm = 101325N/m^2\] \[V = \text{volume of room as m}^3\] Where N is he number of gas particles So the total average thermal energy in the room would be the average kinetic energy multiplied by the number of molecules - then you get an answer reasonable in terms of the order of magnitude. \[\sum K = N K_E = \frac{5}{2}PV\]

OpenStudy (anonymous):

last term should be \[\sum_{i=1}^N \bar{K_i} = N \bar{K} = \frac{5}{2}PV\] where \[\bar{K}_i \overset{i≠j}{=} \bar{K}_j\] Sorry for sloppy conventions :P